OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6 A) -14 B) 24 C) 14 D) -24

OpenStudy (anonymous):

@Pandalover485

OpenStudy (anonymous):

multiply equation 2 by 2 add both and solve for x then find y

OpenStudy (anonymous):

somehow i got A.... I think it is wrong.

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

let me check wait

OpenStudy (anonymous):

2 as in B?

OpenStudy (anonymous):

wait let me cross check no no not B

OpenStudy (anonymous):

okaaaayyyy.. well these are the options i have to choose from.

OpenStudy (anonymous):

i am sorry x=-24 and y=-14

OpenStudy (anonymous):

so i was right, A. right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12 A) 3y + 12 B) -3y + 12 C) 3y - 12 D) -3y - 12

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8 A) (3, 2) B) (2, 3) C) (-2, -3) D) (-2, 3)

OpenStudy (anonymous):

sorry i wrongly deleted comment its still B wait i wil calculate for this 1

OpenStudy (anonymous):

you want method for this 1?

OpenStudy (anonymous):

nahh lol

OpenStudy (anonymous):

um srry

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

you dont? how will you solve in exam o_o

OpenStudy (anonymous):

im in 7th grade but forgot how

OpenStudy (anonymous):

god i wasnt talking to you

OpenStudy (anonymous):

I know how to do this, but I'm stuck on a few, and i wanna get this done as fast as possible >.<

OpenStudy (anonymous):

...this is 9th grade stuff..

OpenStudy (anonymous):

uh huh okay

OpenStudy (anonymous):

so can ya help? C:

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

thanks!!! imma make a new question and tag you in it so i can give you best response again c:

OpenStudy (anonymous):

well i just hope you understand how to solve

OpenStudy (anonymous):

I do(: